Six different national brands of chocolate chip cookies were randomly selected at the supermarket. The grams of fat per serving are as follows: 9; 9; 11; 6; 8; 8. Assume the underlying distribution is approximately normal.

(iii) Calculate the error bound. (Round your answer to two decimal places.)
Answer #1

Solution:

Error bound = Margin of error = E = t*Standard error = t*S/sqrt(n)

From given data, we have

S = 1.643167673

n = 6

df = n – 1 = 5

We assume confidence level = 95%

So, critical t value = 2.5706

(by using t-table)

Standard error = S/sqrt(n)

Standard error = 1.643167673/sqrt(6)

Standard error = 0.67082

Error bound = t*S/sqrt(n)

Error bound = 2.5706*1.643167673/sqrt(6)

Error bound = 2.5706*0.67082

Error bound = 1.72441

Error bound = 1.72